# HP2C-DT: High-Precision High-Performance Computer-enabled Digital Twin

### v1

@@ 6.3 Experiments
In this section, we evaluate the system’s performance through experiments measuring (1) communication bandwidth, (2) response time, and (3) HPC scalability. For this, we use the experimental setup described above. In particular, for (3), we select an application representative of a digital twin of a power system: stability analysis.
@@
However, training such ML models requires data. Although the DT continuously collects measurements from the operation of the actual system, such data may be insufficient to train accurate ML-based surrogate models, which requires the creation of a synthetic dataset [25]. We can generate the synthetic dataset offline by exploring the operational space and applying conventional tools for power system stability assessment at each point. The HP2C-DT framework is well suited for this, as data exploration and generation are highly parallelizable and the framework integrates HPC throughout the digital twin’s architecture. Due to this, we propose a data generator tool to evaluate computation scalability on our prototype.
The tool starts by sampling the entire operational space and progressively focuses on regions of higher interest—those intersecting the stability margin of the system—using an entropy-based criterion [26]. After this, the tool recursively divides the variable space into subregions by segmenting the dimensions with a higher impact on system stability. The execution parameters include the depth of exploration and the branching factor. The variable space consists of the power injected by each generator, the power demanded by each load, and several converter control parameters. The tool is implemented in Python using the GridCal library [27], which specializes in power flow calculations.
@@ 6.3.1 Experiment 1: Reducing communication overhead
We perform the experiment using a single edge node deployed locally, and a cloud node running on an on-premises server at BSC. We assign the edge node a single sensor that generates measurements at predefined acquisition intervals of 1 ms, 10 ms, 100 ms, and 1 s. Each acquisition interval is tested with aggregation intervals of 1 ms, 10 ms, 100 ms, 1 s, and 10 s. The case where both intervals are set to 1 ms represents the baseline, where every measurement is sent individually. The cloud node receives these messages and measures the data transmission rate in bytes per second, including message headers of the RabbitMQ middleware.
@@
Table 2 highlights the impact of actual data aggregation. Since phasor aggregation condenses each rolling window into just two floating-point values (amplitude and angle), the results become independent of the sampling interval. The Required bandwidth decreases proportionally with the aggregation interval, achieving a nearly 10x reduction at each step. In contrast, the all aggregation was limited by the need to transmit every measured value, leading to diminishing improvements.
@@ 6.3.3 Experiment 3: Scalability of HPC workloads
This experiment evaluates how well HPC resources scale when handling computationally intensive tasks. As a test case, we use the power grid data generator described earlier, applied to the IEEE 118-bus transmission model [28]. The experiment explores a variable space that includes the power delivered by the 53 converter, the proportion of grid-following vs. grid-forming operation for each converter, the power consumed by the 91 loads, and the control parameters of each converter. The computations are parallelized and distributed using the COMPSs framework and executed on the MareNostrum 5 supercomputer at BSC.
@@
Beyond raw scalability, the practical impact of HPC resources is significant. With 32 nodes, the execution time is reduced to approximately one hour, compared to multiple hours on a single high-performance server. Even with 16 nodes, the execution takes below two hours. In a digital twin scenario, this level of performance would enable real-time grid stability assessments with hourly updates, as opposed to daily recalculations. This higher update frequency would significantly improve prediction accuracy and system responsiveness, which would be impractical in a cloud-based infrastructure due to cost and inconsistent performance.
@@ 7 Conclusion
Future work will focus on decentralizing the architecture by organizing edge nodes into subclusters to improve availability and enable fail-resilient configurations. Additionally, efforts will focus on refining the lifecycle of digital twin models—whether physics-based or data-driven—to ensure their continuous adaptation and deployment across Edge and Cloud layers. This includes integrating adaptive machine learning capabilities that autonomously learn from sensed data, allowing real-time refinement of the digital twin.

### v2

@@ 6.3 Experiments
In this section, we evaluate the system’s performance through experiments that measure (1) communication bandwidth, (2) response time, and (3) HPC scalability. For this, we use the experimental setup described above. In particular, for (3), we select an application representative of a digital twin of a power system: stability analysis, and a standard case study, the IEEE 118-bus transmission model [27].
@@
However, training such ML models requires data. Although the digital twin continuously collects measurements from the operation of the actual system, such data may be insufficient to train accurate ML-based surrogate models, which requires the creation of a synthetic dataset [29]. We can generate the synthetic dataset offline by exploring the operational space and applying conventional tools for power system stability assessment at each point. The HP2C-DT framework is well suited for this, as data exploration and generation are highly parallelizable and the framework integrates HPC throughout the digital twin’s architecture. Due to this, we propose a data generator tool to evaluate computation scalability on our prototype.
As shown in Algorithm 1, the exploration proceeds by sampling the entire operational space and progressively focuses on regions of higher interest—those intersecting the stability margin of the system—using an entropy-based criterion [30]. After this, the tool recursively divides the variable space into subregions by segmenting the dimensions with a higher impact on system stability. Execution parameters include search space range, number of samples per subregion, and depth of exploration. The variable space consists of the power injected by each generator, the power demanded by each load, and several converter control parameters. The tool is implemented in Python using the VeraGrid library (previously known as GridCal [31]), for feasible operating point power flow calculation, and the Python version of STAMP [32], for an EMT-focused small-signal stability analysis.
Algorithm 1 Entropy-based Recursive Subregion Exploration
1:Dimensions of the search space and their ranges, number of samples per cell N, objective function Eval_Stability, maximum depth d, grid configuration, use_sensitivity
2:Generated dataset with power flow results and stability labels
3:root_cell $\leftarrow$ entire exploration space
4:depth $\leftarrow 0$
5:entropy_parent $\leftarrow 0$
6:dataset $\leftarrow$ empty
7:Explore_Cell(root_cell, depth, entropy_parent)
8:function Explore_Cell(cell, depth, entropy_parent)
9:  samples $\leftarrow$ Latin_Hypercube(cell, N)
10:  results $\leftarrow$ empty
11:  for all sample in samples do parallel
12:   result $\leftarrow$ Eval_Stability(sample, grid configuration)
13:   results.append(result)
14:  end for
15:  dataset.append(results)
16:  entropy_child $\leftarrow$ Eval_Entropy(results)
17:  if entropy_child >> entropy_parent and depth << d then
18:   subcells $\leftarrow$ Divide_Cell(cell, use_sensitivity)
19:   for all subcell in subcells do parallel
20:     Explore_Cell(subcell, depth + 1, entropy_child)
21:   end for
22:  end if
23:end function
24:function Eval_Entropy(samples)
25:  p_stable $\leftarrow$ fraction of samples labeled stable
26:  p_unstable $\leftarrow$ 1 - p_stable
27:  return $(\text{p\_stable}\log_{2}\text{p\_stable}+\text{p\_unstable}\log_{2}\text{p\_unstable})$
28:end function
29:function Divide_Cell(cell, use_sensitivity)
30:  if use_sensitivity then
31:   Divide along the most sensitive dimension(s)
32:  else
33:   Divide along preconfigured grid dimensions
34:  end if
35:  return list of subcells
36:end function
@@ 6.3.1 Experiment 1: Reducing communication overhead
We perform the experiment using a single edge node deployed locally and a cloud node running on an on-premises server at BSC. We assign the edge node a single sensor that generates measurements at predefined acquisition intervals of 1 ms, 10 ms, 100 ms, and 1 s. Each acquisition interval is tested with aggregation intervals of 1 ms, 10 ms, 100 ms, 1 s, and 10 s. The case where both intervals are set to 1 ms represents the baseline, where every measurement is sent individually. The cloud node receives these messages and measures the data transmission rate in bytes per second, including message headers of the RabbitMQ middleware.
@@
Figure 4: Experiment 1: Required bandwidth vs. sampling interval $T_{s}$ for varying numbers of edge nodes ($N=1,2,4,8$) at fixed aggregation interval $T_{a}=10$ s. Subfigure (a) shows the all aggregator, where bandwidth decreases with $T_{s}$ and scales linearly with the number of edges. Subfigure (b) shows the phasor aggregator, where bandwidth remains independent of $T_{s}$ but scales linearly with the number of edges.
Table 2 highlights the impact of actual data aggregation. Since phasor aggregation condenses each rolling window into just two floating-point values (amplitude and angle), the results become independent of the sampling interval. The Required bandwidth decreases proportionally with the aggregation interval, achieving a nearly 10x reduction at each step. In contrast, the all aggregation is limited by the need to transmit every measured value, leading to diminishing improvements.
Figure 5 more explicitly compares the effect of using the two extreme aggregators and shows how the compression gain evolves with the sampling interval. Compression gain is defined as the ratio of bandwidth required by all aggregation to that required by phasor aggregation (see Equation 7). At very fine sampling ($T_{s}=1$ ms), the phasor aggregator condenses a large number of measurements per aggregation window, resulting in a massive compression ($\approx$866). As $T_{s}$ increases, fewer raw measurements fit inside each window, and the benefit decreases to $\approx$1.8 for $T_{s}=1000$ ms. Compression is most effective when the ratio $T_{s}/T_{a}$ is high and fades as $T_{s}$ and $T_{a}$ get closer.
$\text{Compression Gain}=\frac{B_{\mathrm{all}}}{B_{\mathrm{phasor}}}$ (7)
Figure 5: Experiment 1: Compression gain comparison for $T_{a}=10$ s.
We now extend the experiment to multiple edge nodes to assess the impact of scaling on bandwidth requirements. Figure 4 shows how the required bandwidth scales with the number of edge nodes when varying the sampling interval $T_{s}$, fixing the aggregation interval at $T_{a}=10$ s. This choice ensures the widest coverage along the $x$-axis (since there are no samples for $T_{s}>T_{a}$) and also highlights the maximum effect of aggregation, as larger windows provide the strongest bandwidth reduction. Subfigure (a) confirms that for the all aggregator, the bandwidth decreases with $T_{s}$ following the same trend across all scales. The four curves are essentially parallel and evenly spaced in logarithmic scale, reflecting a nearly ideal linear scaling with the number of edges at small to moderate scales. Subfigure (b) shows the case for the phasor aggregator, where the results are independent of $T_{s}$ and the curves flatten into horizontal lines. As expected, the bandwidth requirements simply scale with the number of edge nodes, but remain constant across different sampling rates owing to the strong compression inherent to the phasor method. Together, these results show that the observed trends in single-node experiments extend to multi-node setups.
@@
While this experiment focuses on measuring bandwidth reduction, it also sets the stage for evaluating trade-offs between communication overhead, data freshness, and accuracy. In this specific setup, both the all and phasor aggregators preserve the signal, so no loss is observed. Still, the framework is flexible and allows domain-specific aggregators and real-world signals, where some information loss may occur. The metrics and approach we used here can easily be extended to study reconstruction error or data age, providing a solid basis for future experiments on these trade-offs in edge–cloud digital twins.
@@ 6.3.3 Experiment 3: Scalability of HPC workloads
This experiment evaluates how well HPC resources scale when handling computationally intensive tasks. The experiment explores a variable space that includes the power delivered by the 53 converter, the proportion of grid-following vs. grid-forming operation for each converter, the power consumed by the 91 loads, and the control parameters of each converter. The computations are parallelized and distributed using the COMPSs framework and executed on the MareNostrum 5 supercomputer at BSC.
@@
Beyond raw scalability, the practical impact of HPC resources is significant. With 32 nodes, the execution time is reduced to approximately one hour, compared to multiple hours on a single high-performance server. Even with 16 nodes, the execution takes below two hours.
In a digital twin scenario, this level of performance enables high-volume offline data generation to be completed within approximately one hour rather than several days. Such accelerated data production allows operators to frequently update or retrain machine learning models for real-time stability assessment, or to refresh stability-related constraints used in real-time optimization. This capability enhances prediction accuracy and system adaptability by keeping models aligned with evolving grid conditions. Achieving this level of computational throughput would be impractical in a purely cloud-based infrastructure due to cost and performance variability.
7 Discussion
The experiments presented in this work focus on different aspects of the HP2C-DT framework, including data aggregation (Experiment 1), function execution and offloading (Experiment 2), and HPC scalability (Experiment 3). Although a complete system-level benchmark would provide a complete picture of end-to-end performance, such evaluations are highly domain- and infrastructure-dependent and beyond the scope of this study. Still, these experiments provide useful information on the behavior of the framework, help clarify infrastructure needs for domain-specific applications (particularly Experiment 1), and highlight its potential for integrated operation across Edge, Cloud, and HPC layers.
Experiment 1 highlights the advantages of targeted aggregation techniques for reducing communication, while also showing the overhead introduced by the messaging middleware. Regarding messaging overhead, RabbitMQ was chosen not only for raw performance, but for its maturity, reliability, ease of integration with containerized deployments, and native support for publish-subscribe semantics. These properties were important during early architectural development to ensure robustness and flexibility. Rolling-window mechanisms complement this setup, not as a workaround for RabbitMQ’s overhead, but as a strategy to manage bandwidth and scalability across the edge-cloud boundary; something that would likely be needed even with a leaner messaging backend.
Experiment 2 focuses on function execution, comparing sequential processing on the edge, parallel workflows on the edge, and fully or partially offloading to the cloud. The results show that offloading can reduce response times for larger computational tasks and that the framework automatically adapts to available resources while supporting synchronous and asynchronous execution. This illustrates how the architecture can leverage cloud resources without disrupting edge operations.
Experiment 3 evaluates HPC scalability, demonstrating how COMPSs not only supports intra-node parallelization and cloud offloading but also enables substantial speedups for computationally intensive tasks. Even with very large workloads, performance scales well up to 32 nodes (3,584 processors), reducing execution times from several hours to roughly one hour. In the full system, the cloud acts as the main coordination layer, executing decision-making logic, and running mid-tier simulations that complement fast edge-level updates. The HPC cluster is exposed to the cloud as an available resource and is typically used for heavy asynchronous computation campaigns, such as less frequent model updates or broader operational space exploration. This division of roles is what allows the framework to balance responsiveness with computational depth, with the cloud managing the flow of tasks and the HPC layer absorbing the most demanding workloads. Improvements in performance at the HPC layer allow for more frequent updates to digital twin models, which in turn enhances prediction accuracy and system responsiveness of the entire edge-cloud-HPC architecture.
Due to practical reasons, Python is included in HP2C-DT to facilitate the development of digital twins. Its versatility and the ecosystem of AI and data analysis libraries surrounding it make it valuable for modeling and experimentation. Java, on the other hand, smooths the integration of COMPSs and takes care of the time-critical parts of the framework, such as scheduling, orchestration, and hardware interfacing. Java provides runtime-level portability and isolation through the JVM, while Docker containers add system-level packaging (network, file system, process space, dependencies) and make deployment consistent across heterogeneous nodes. This division of roles, together with containerization, ensures that developers can prototype and integrate domain-specific models with minimal friction, without sacrificing execution efficiency where it matters.
Taken together, these observations suggest that HP2C-DT provides a flexible and robust foundation for edge–cloud–HPC digital twin applications. Each subsystem experiment contributes to understanding the behavior of the framework, and the architecture is easily extensible to domain-specific tasks and metrics in future evaluations.
@@ 8 Conclusion
Future work will focus on decentralizing the architecture by organizing edge nodes into subclusters to improve availability and enable fail-resilient configurations. Additionally, efforts will focus on refining the lifecycle of digital twin models—whether physics-based or data-driven—to ensure their continuous adaptation and deployment across Edge and Cloud layers. This includes further testing how Cloud and HPC resources can be combined to trade Cloud flexibility against HPC queue-based computational power, and integrating adaptive machine learning capabilities that autonomously learn from sensed data, allowing real-time refinement of the digital twin.